OpenStudy (hitaro9):

Let C be parametrized by x = 10 sin^2(t) and y = 10 sin (t) cos (t) for 0 ≤ t ≤ π/2 Find the length L of C.

OpenStudy (lastdaywork):

cos(2t) = 5 - x sint(2t) = y Hence, this is a circle. Try to find the length of arc C (using mental maths)

OpenStudy (hitaro9):

How'd you get those two exactly?

OpenStudy (lastdaywork):

Use trigonometry to reduce the powers ; see formulas \[\cos(2t) = 1 - 2\sin^2(t)\] \[\sin(2t) = 2\sin(t) \cos(t)\]